OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = -3 over 4x - 2? y = 4 over 3x - 12 y = 3 over 4x - 5 y = -4 over 3x + 20 y = -3 over 4x + 13

OpenStudy (haseeb96):

first take out the slope of from the given equation y=mx+b always constant term in this type of equation is yur slope so here what will be the slope it is yur job to find the slope here

OpenStudy (anonymous):

-3/4

OpenStudy (haseeb96):

and reverse it because x is in the denominator so it will become m=-4/3

OpenStudy (anonymous):

so its C

OpenStudy (haseeb96):

our work is not done still ... @kellyconnelly

OpenStudy (anonymous):

y = -4/3x + 20

OpenStudy (haseeb96):

here the line is perpendicular so use this condition of slope here m1m2=-1 and m1=-4/3 (-4/3)m2=-1 m2 = 3/4

OpenStudy (haseeb96):

now u can answer the question logically that which one is correct

OpenStudy (haseeb96):

B is the correct answer

OpenStudy (haseeb96):

now use here the point slope form and get the answer which is B

OpenStudy (driftracer305):

k @kellyconnelly is it \[y = \frac{ -3 }{ 4x-2 } ......... or............ y = \frac{ -3 }{ 4x } -2....? \]

OpenStudy (driftracer305):

omg......... which one...... the first one or the second one??

OpenStudy (haseeb96):

if it is second one then A is the correct answer

OpenStudy (haseeb96):

@kellyconnelly tell me how the equation is written in yur book

OpenStudy (anonymous):

its neither

OpenStudy (anonymous):

its 3/4 and then an x

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }x\] but negative

OpenStudy (anonymous):

@Haseeb96 @driftracer305

OpenStudy (haseeb96):

so A is the correct answer

OpenStudy (haseeb96):

@kellyconnelly just revise the method again because in the previous i thought that it is -3/4x and x is the denomenator but i was wrong it is in the numenator

OpenStudy (haseeb96):

so here we have no need to reverse it just put the slope into the perpendicular condition and get the answer

OpenStudy (anonymous):

\[y = -\frac{ 3 }{ 4 }x - 2\]